OpenStudy (anonymous):

What would be a graph for y = x2? I'm stumped.

OpenStudy (anonymous):

for this? \[y=x^2\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I will help you with a couple of points, but it's a curve, not a line, so don't just "connect the dots"

OpenStudy (anonymous):

The curve shape is called a parabola

OpenStudy (anonymous):

would it be a curve like a U

OpenStudy (anonymous):

yes, mostly. It happens to be centered on the origin, (0,0), because when x is 0, y = 0^2 = 0

OpenStudy (anonymous):

OH! I get it :)

OpenStudy (anonymous):

For every positive x like x = 2, y = x^2 = 2^2 = 4 But for the mirror image negative x's, like x = -2, y also equals 4 The graph has symmetry around the y axis

OpenStudy (anonymous):

Great! glad to help :)